Algebra

Find number of permutation and combination of four letters taken away from the word EXAMINATION.

Pratul kumar
12 Years agoGrade

1 Answer

Sher Mohammad
12 Years ago
A, I , N occurs twice
total distinct letters=8
selecting 4 letters all distinct=8C4=70
Selecting 2 same letters and 2 distinct =3C1*7C2=63
selecting 2same, 2same =3C2=3
total=136

and permutation, 8C4*4!+3C1*7C2*4!/2!+3C2*4!/2!2! = 24*70+63*12+3*6=2454

Sher Mohammad
B.Tech, IIT Delhi.